# fourierlargesymbols147 thanks: Present address for Q. H. Tran: Quantum Laboratory, Fujitsu Research, Fujitsu Limited

### v2

@@ II.1 Non-stationary ESP
$\mathrm{Var}_{w}^{t}(\{\mathbf{u}_{\tau}\})$ $\equiv{\mathrm{E}}_{w}^{t}\left(\{\|\mathbf{u}_{\tau}-\overline{\mathrm{E}}_{w}^{t}(\{\mathbf{u}_{\tau}\})\|^{2}\}\right)$
@@ II.3 Non-stationary ESP of QRC
$U(\mathbf{u};\theta)$ $\equiv U_{3}^{\dagger}(\theta)R_{Z}(\mathbf{u})U_{3}(\theta)$ (9)
@@ II.3.1 ESP and non-stationary ESP
First, we computed the ESP indicator, the left-hand side of Eq. (1), and non-stationary ESP indicator, the left-hand side of Eq. (2), for every configuration of the parameter $\theta$. We randomly sampled four input sequences of size $200$ from $\mathrm{Uniform}([-1,1])$ and three initial reservoir states from Haar random distribution of 2-qubit pure states.
@@ II.3.3 MC and IPC
$v_{t}=\prod_{i}Y_{d_{i}}(u_{t-k_{i}})\quad\text{s.t.}\quad\sum_{i}=d_{i}.$ (16)
@@ II.4 Subset non-stationary ESP of QRC
The input encoding also has a tensor product structure. Namely, given an input $u\in[-1,1]$, our input encoding unitary $U_{in}$ was
$U_{in}(u)\equiv R_{Y}^{(0)}(u\pi)\otimes R_{Y}^{(1)}(u\pi).$ (21)
@@
$\rho_{t+1}$ $=\mathcal{E}_{in}\left(\mathcal{E}_{sys}\left(\rho\right),u_{t}\right)\ \ \ \textrm{where}$ (22)
$\mathcal{E}_{in}(\rho,u_{t})\equiv U_{in}(u)\rho U_{in}^{\dagger}(u).$
@@ II.4.1 Subset ESP and subset non-stationary ESP
Similar to the setup in the non-stationary ESP experiments, we computed the non-stationary ESP indicator (the left-hand side of Eq. (2)) for every configuration of $\gamma$ and $p$. The other setups are the same in the non-stationary ESP experiment in Sec II.3.1. We randomly sampled four input sequences of size $200$ from a uniform distribution over the interval $[-1,1]$ and three initial reservoir states from Haar random distribution of 2-qubit pure states.
@@ Appendix A Equivalence of ESP definitions
The following ESP conditions are equivalent. Following the notations introduced in [27], let us consider input sequences $\{\mathbf{u}_{t}\}_{t\in\mathcal{T}}\in U^{\mathcal{T}}$ where $U$ is compact. Let $U^{-\mathbb{N}}$ be a left-infinite sequence up to some $t\geq 0$, namely, $U^{-\mathbb{N}}=\{\cdots,\mathbf{u}_{t-1},\mathbf{u}_{t}\}$.
@@
    We prove the contraposition of the fact. Namely, if $\mathbf{x}_{t}$ is not an echo function, then there exists an input sequence $\{\mathbf{u}_{t}\}$ such that $\mathbf{x}_{t}\neq\mathbf{x}_{t}^{\prime}$. Indeed, if $\mathbf{x}_{t}$ is not an echo function, there exists at least one input-independent parameter $\theta$ such that $\mathbf{x}_{t}=\mathbf{E}^{\prime}(\{\mathbf{u}_{t}\},\theta)$. This form implies that the effect of $\theta$ remains finite after processing an infinite number of inputs for at least one input sequence because $\{\mathbf{u}_{t}\}$ is left-infinite. Therefore, for different states $\mathbf{x}_{t}$ and $\mathbf{x}_{t}^{\prime}$ with different $\theta$, there exists at least one input sequence $\{\mathbf{u}_{t}\}\in U^{-\mathbb{N}}$ such that $\mathbf{x}_{t}\neq\mathbf{x}_{t}^{\prime}$. This proves the sufficiency of the remark.
@@
One of the confusions about the equality of the conditions comes when we let $\theta=t$, a time parameter. Since input sequence $\{\mathbf{u}_{t}\}$ also has time parameter $t$, we are keen to equate them. However, the time parameter $t$ of $\{\mathbf{u}_{t}\}$ is a constant between all dynamics with different initial states for each time step: a function of the input sequence itself. Namely, for states $\mathbf{x}$ and $\mathbf{x}^{\prime}$ having the same input sequence $\{\mathbf{u}_{t}\}$ at every input cycle, $t$, the input cycle, does not differentiate their states. Therefore, if we set a time parameter $\theta=\tau$ as a parameter of the echo function, $\tau$ must not depend on the timestamp of the input sequence. It is an unrelated parameter to the cycle of the data input procedure and typically an initial time $t_{0}$. There is a work [29] in which temporal information processing capacity (TIPC) analyzes the input cycle-dependent structure of state sequences. We argue that all input sequences are input cycle-dependent. TIPC examines how initial state dependency, which does not converge after washout, evolves through the input cycle. Namely, it analyzes the dynamics of initial state dependency.

### v3

@@ II.1 Non-stationary ESP
$\mathrm{Var}_{w}^{t}(\{\mathbf{u}_{\tau}\})$ $\equiv{\mathrm{E}}_{w}^{t}\left(\{\|\mathbf{u}_{\tau}-\mathrm{E}_{w}^{t}(\{\mathbf{u}_{\tau}\})\|^{2}\}\right)$
@@ II.3 Non-stationary ESP of QRC
$U(\mathbf{u};\theta)$ $\equiv U_{3}^{\dagger}(\theta)R_{Z}(\mathrm{arccos}(\mathbf{u}))U_{3}(\theta)$ (9)
@@ II.3.1 ESP and non-stationary ESP
First, we computed the ESP indicator, defined as follows:
$\mathcal{I}_{ESP}(t,s_{0},s_{0}^{\prime})\equiv\frac{\|f(\{\mathbf{u}_{\tau}\}_{\tau\leq t};s_{0})-f(\{\mathbf{u}_{\tau}\}_{\tau\leq t};s_{0}^{\prime})\|}{\|s_{0}-s_{0}^{\prime}\|},$ (12)
and the non-stationary ESP indicator defined as follows:
$\mathcal{I}_{NS}(t,s_{0},s_{0}^{\prime},w)\equiv$ (13)
$\quad\mathcal{I}_{ESP}(t,s_{0},s_{0}^{\prime})\times\frac{\sqrt{\min\left[\overline{\mathrm{Var}}_{w}^{w}(f;s_{0}),\overline{\mathrm{Var}}_{w}^{w}(f;s^{\prime}_{0})\right]}}{\sqrt{\min\left[\overline{\mathrm{Var}}_{w}^{t}(f;s_{0}),\overline{\mathrm{Var}}_{w}^{t}(f;s^{\prime}_{0})\right]}}$
for every configuration of the parameter $\theta$. We randomly sampled four input sequences of size $200$ from $\mathrm{Uniform}([-1,1])$ and three initial reservoir states from Haar random distribution of 2-qubit pure states.
@@ II.3.3 MC and IPC
$v_{t}=\prod_{i}Y_{d_{i}}(u_{t-k_{i}})\quad\text{s.t.}\quad\sum_{i}d_{i}=d.$ (18)
@@ II.4 Subset non-stationary ESP of QRC
The input encoding also has a tensor product structure. Namely, given an input $\mathbf{u}\in[-1,1]$, our input encoding unitary $U_{in}$ was
$U_{in}(u)\equiv R_{Y}^{(0)}(\mathrm{arccos}(\mathbf{u}))\otimes R_{Y}^{(1)}(\mathrm{arccos}(\mathbf{u})).$ (23)
@@
$\rho_{t+1}$ $=\mathcal{E}_{in}\left(\mathcal{E}_{sys}\left(\rho\right),\mathbf{u}_{t}\right)\ \ \ \textrm{where}$ (24)
$\mathcal{E}_{in}(\rho,\mathbf{u}_{t})\equiv U_{in}(\mathbf{u}_{t})\rho U_{in}^{\dagger}(\mathbf{u}_{t}).$
@@ II.4.1 Subset ESP and subset non-stationary ESP
Similar to the setup in the non-stationary ESP experiments, we computed the non-stationary ESP indicator (Eq. (13)) for every configuration of $\gamma$ and $p$. The other setups are the same in the non-stationary ESP experiment in Sec II.3.1. We randomly sampled four input sequences of size $200$ from a uniform distribution over the interval $[-1,1]$ and three initial reservoir states from Haar random distribution of 2-qubit pure states.
@@ Appendix A Equivalence of ESP definitions
The following ESP conditions are equivalent.
Following the notations introduced in [27], let us consider input sequences $\{\mathbf{u}_{t}\}_{t\in\mathcal{T}}\in U^{\mathcal{T}}$ where $U$ is compact. Let $U^{-\mathbb{N}}$ be a left-infinite sequence up to some $t\geq 0$, namely, $U^{-\mathbb{N}}=\{\cdots,\mathbf{u}_{t-1},\mathbf{u}_{t}\}$.
@@
    We prove the contraposition of the fact. Namely, if $\mathbf{x}_{t}$ is not an echo function, then there exists an input sequence $\{\mathbf{u}_{t}\}$ such that $\mathbf{x}_{t}\neq\mathbf{x}_{t}^{\prime}$.
    Indeed, if $\mathbf{x}_{t}$ is not an echo function, there exists at least one input-independent parameter $\theta$ such that $\mathbf{x}_{t}=\mathbf{E}^{\prime}(\{\mathbf{u}_{t}\},\theta)$. This form implies that the effect of $\theta$ remains finite after processing an infinite number of inputs for at least one input sequence because $\{\mathbf{u}_{t}\}$ is left-infinite. Therefore, for different states $\mathbf{x}_{t}$ and $\mathbf{x}_{t}^{\prime}$ with different $\theta$, there exists at least one input sequence $\{\mathbf{u}_{t}\}\in U^{-\mathbb{N}}$ such that $\mathbf{x}_{t}\neq\mathbf{x}_{t}^{\prime}$. This proves the sufficiency of the remark.
@@
One of the confusions about the equality of the conditions comes when we let $\theta=t$, a time parameter. Since input sequence $\{\mathbf{u}_{t}\}$ also has time parameter $t$, we are keen to equate them. However, the time parameter $t$ of $\{\mathbf{u}_{t}\}$ is a constant between all dynamics with different initial states for each time step: a function of the input sequence itself. Namely, for states $\mathbf{x}$ and $\mathbf{x}^{\prime}$ having the same input sequence $\{\mathbf{u}_{t}\}$ at every input cycle, $t$, the input cycle, does not differentiate their states.
Therefore, if we set a time parameter $\theta=\tau$ as a parameter of the echo function, $\tau$ must not depend on the timestamp of the input sequence. It is an unrelated parameter to the cycle of the data input procedure and typically an initial time $t_{0}$.
There is a work [29] in which temporal information processing capacity (TIPC) analyzes the input cycle-dependent structure of state sequences. We argue that all input sequences are input cycle-dependent. TIPC examines how initial state dependency, which does not converge after washout, evolves through the input cycle. Namely, it analyzes the dynamics of initial state dependency.